# gRNAde: Geometric Deep Learning for 3D RNA inverse design

### v5

@@ 4.2 Multi-state RNA design benchmark
Multi-state gRNAde boosts sequence recovery. In Figure 6(a), we compared a single-state variant of gRNAde with otherwise equivalent multi-state models (with up to 3 and 5 states, respectively) in terms of native sequence recovery  [^3] Self-consistency scores for multi-state design are perhaps less reliable than single-state design due to the lack of multi-state structure prediction tools. At present, we compute the average score between multiple ground truth structures and the predicted structure for designed sequences.. Multi-state variants show marginal improvements, overall. As a caveat, it is worth noting that multi-state models consume more GPU memory than an equivalent single-state model during mini-batch training (approximate peak GPU usage for max. number of states = 1: 12GB, 3: 28GB, 5: 50GB on a single A100 with at most 3000 total nodes in a mini-batch).
Improved recovery in structurally flexible regions. In Figure 6(b), we evaluated gRNAde’s multi-state sequence recovery at a fine-grained, per-nucleotide level. Multi-state GNNs improve sequence recovery over the single-state variant on structurally flexible nucleotides, as characterised by undergoing changes in base pairing/secondary structure, higher average RMSD between 3D coordinates across states, and larger solvent accessible surface area.
(a) Per-sample sequence recovery
(b) Per-nucleotide recovery vs. structural flexibility measures
Figure 6: Multi-state design benchmark. (a) Multi-state gRNAde show marginal improvement over an equivalent single-state model in terms of average per-sample sequence recovery over all test RNAs. (b) When plotting sequence recovery per-nucleotide, multi-state gRNAde improves over a single-state model for structurally flexible regions of RNAs, as characterised by nucleotides that tend to undergo changes in base pairing (left), nucleotides with greater average solvent accessible surface area (centre), and nucleotides with higher average RMSD (right) across multiple states. Marginal histograms in blue show the distribution of values. We plot performance for one consistent random seed across all models; collated results and ablations are available in Appendix Table 1.
@@ 4.3 Zero-shot ranking of RNA fitness landscape
gRNAde outperforms random baselines in low design budget scenarios. Figure 7 illustrates the results of our retrospective study. At low design budgets of up to hundreds of sequences, which are relevant in the case of a low throughput fitness screening assay, gRNAde outperforms all random baselines in terms of the maximum change in fitness over the wild type. The top 10 mutants as ranked by gRNAde contain a sequence with 4-fold improved fitness, while the top 200 leads to a 5-fold improvement555 As a caveat, the fitness assays from McRae et al. (2024) used for creating the landscape have inherent noise and cannot easily differentiate between mutants of similar activity.. Note that gRNAde is used zero-shot here, i.e. it was not fine-tuned on any assay data.
Perspective. Overall, it is promising that gRNAde’s perplexity correlates with experimental fitness measurements out-of-the-box (zero-shot) and can be a useful ranker of mutant fitness in our retrospective study. In realistic design scenarios, improvements could likely be obtained by fine-tuning gRNAde on a low amount of experimental fitness data. For example, latent features from gRNAde may be finetuned or used as input to a prediction head with supervised learning on fitness landscape data. This study acts as a sanity check before committing to wet lab validation of gRNAde designs. We see random mutagenesis and directed evolution-based approaches as complementary to de-novo design and inverse folding approaches like gRNAde. Random mutagenesis can be thought of as local exploration around a wild type sequence, optimising fitness within an ‘island’ of activity. Structure-based design approaches are akin to global jumps in sequence space, with the potential to find new islands further away from the wild type (Huang et al., 2016).
@@ 5 Conclusion
We introduce gRNAde, a geometric deep learning pipeline for RNA sequence design conditioned on one or more 3D backbone structures. gRNAde is superior to the physically based Rosetta for 3D RNA inverse folding in terms of performance, inference speed, and ease of use. Further, gRNAde enables explicit multi-state design for structurally flexible RNAs which was previously not possible with Rosetta. gRNAde’s perplexity correlates with native sequence and structural recovery, and can be used for zero-shot ranking of mutants in RNA engineering campaigns. To the best of our knowledge, gRNAde is also the first geometric deep learning architecture for multi-state biomolecule representation learning; the model is generic and can be repurposed for other learning tasks on conformational ensembles, including multi-state protein design.
Key avenues for future development of gRNAde include supporting multiple interacting chains, accounting for partner molecules with RNAs, and supporting negative design against undesired conformations. We discuss practical tradeoffs to using gRNAde in real-world RNA design scenarios in Appendix C, including limitations due to the current state of 3D RNA structure prediction tools. Finally, we are hopeful that advances in RNA structure determination and computationally assisted cryo-EM (Kappel et al., 2020; Bonilla and Kieft, 2022) will further increase the amount of RNA structures available for training geometric deep learning models in the future.
@@ Acknowledgements
Appendix A 3D Visualisation of gRNAde Designs
@@
Figure 9: Cherry-picked designs for Guanine riboswitch aptamer (PDB: 4FE5, sequence: GGACAUAUAAUCGCGUGGAUAUGGCACGCAAGUUUCUACCGGGCACCGUAAAUGUCCGACUAUGUCC).
@@
GGGGCUCCGGCGACGCAGUCGAAAG
CCCAGCAGUACCAAGCCUCAGGGGA
AACUUUGAGGUGGCCUAACAAAGGA
UACGGUAAUAAGCUGCGGGAAAAGG
UUGUAAGCCGGAGCGAAGACCUAAG
GCACCGCUUUUGGCGGUGCUAUGGU
UGAAGUUAA
@@
GGGGUACCGGCGACGCAGUCGAAUG
CCCUGUGGUACCAAGCCCCGGGGGA
AACUUCGGGGUGGCCUUACCAAGGA
CACGGUAAUAAGCCACGGGAAAUGG
UUGUAAGCCGGUCCGAAGCCCUAAG
GCCGCGCUUUGGGCGCGGCUAUGGG
UGAAGGCAA
@@
GAGGCCACGGCAACGCAGUCUAACG
CCCUGUGGUACCAAGUCUUAGGAGA
AAUUUUAAGAUGGCCUAAUAAAGGA
UAUGGUAAUAAGCCACGGGAAAAGG
UUGUAAGACGUGACGAAGUCCUAAG
GCCACAGUUUUGCUGUGGCUAUGGA
UGGAGUACA
@@
Figure 10: Cherry-picked designs for Tetrahymena Ribozyme P4-P6 domain (PDB: 2R8S, sequence: GGAAUUGCGGGAAAGGGGUCAACAGCCGUUCAGUACCAAGUCUCAGGGGAAACUUUGAGAUGGCCUUGCAAAGGGU
AUGGUAAUAAGCUGACGGACAUGGUCCUAACACGCAGCCAAGUCCUAAGUCAACAGAUCUUCUGUUGAUAUGGAUGCAGUUCA

### v6

@@ 4.2 Multi-state RNA design benchmark
Multi-state gRNAde consistently boosts sequence recovery. In Figure 5(a), we compared a single-state variant of gRNAde with otherwise equivalent multi-state models (with up to 5 states) in terms of native sequence recovery. Multi-state variants show a consistent 3-5% improvement, with the best performance obtained using 3 states. This trend holds to a lesser extent on the single-state benchmark where the multi-state model is being used with only one state as input. This suggests that seeing multiple states during training can be useful for teaching gRNAde about RNA conformational flexibility and improve performance even for single-state design tasks. As a caveat, it is worth noting that multi-state models consume more GPU memory than an equivalent single-state model during mini-batch training (approximate peak GPU usage for max. number of states = 1: 12GB, 3: 28GB, 5: 50GB on a single A100 with at most 3000 total nodes in a mini-batch).
Improved recovery in structurally flexible regions. In Figure 5(b), we evaluated gRNAde’s multi-state sequence recovery at a fine-grained, per-nucleotide level to understand the source of performance gains. Multi-state GNNs improve sequence recovery over the single-state variant on structurally flexible nucleotides, as characterised by undergoing changes in base pairing/secondary structure and higher average RMSD between 3D coordinates across states.
@@ 4.3 Zero-shot ranking of RNA fitness landscape
gRNAde outperforms random baselines in low design budget scenarios. Figure 6 illustrates the results of our retrospective study. At low design budgets of up to hundreds of sequences, which are relevant in the case of a low throughput fitness screening assay, gRNAde outperforms all random baselines in terms of the maximum change in fitness over the wild type. The top 10 mutants as ranked by gRNAde contain a sequence with 4-fold improved fitness, while the top 200 leads to a 5-fold improvement. Note that gRNAde is used zero-shot here, i.e. it was not fine-tuned on any assay data.
Overall, it is promising that gRNAde’s perplexity correlates with experimental fitness measurements out-of-the-box (zero-shot) and can be a useful ranker of mutant fitness in our retrospective study. In realistic design scenarios, improvements could likely be obtained by fine-tuning gRNAde on a low amount of experimental fitness data. For example, latent features from gRNAde may be finetuned or used as input to a prediction head with supervised learning on fitness landscape data.
@@ 5 Conclusion
We introduce gRNAde, a geometric deep learning pipeline for RNA sequence design conditioned on one or more 3D backbone structures. gRNAde represents a significant advance over physics-based Rosetta in terms of both computational and experimental performance, as well as inference speed and ease-of-use. Further, gRNAde enables explicit multi-state design for structurally flexible RNAs which was previously not possible with Rosetta. gRNAde’s perplexity correlates with native sequence and structural recovery, and can be used for zero-shot ranking of mutants in RNA engineering campaigns. gRNAde is also the first geometric deep learning architecture for multi-state biomolecule representation learning; the model is generic and can be repurposed for other learning tasks on conformational ensembles, including multi-state protein design.
Limitations. Key avenues for future development of gRNAde include supporting multiple interacting chains, accounting for partner molecules with RNAs, and supporting negative design against undesired conformations. We discuss practical tradeoffs to using gRNAde in real-world RNA design scenarios in Appendix F, including limitations due to the current state of 3D RNA structure prediction tools.
@@ Acknowledgements
Appendices
[…]
[appendix]1
Appendix A Related Work
Appendix B Ablation Study
[…]
Self-consistency metrics
Max. Max. train Perplexity Native seq. 2D – EternaFold 3D – RhoFold
Split #states Model GNN length ($\downarrow$) recovery ($\uparrow$) scMCC ($\uparrow$) scRMSD ($\downarrow$) scTM-score ($\uparrow$) scGDT_TS ($\uparrow$)
Single-state split 1 AR Equiv 500 1.77$\pm$0.07 0.438$\pm$0.01 0.624$\pm$0.07 13.01$\pm$1.18 (0.5%) 0.21$\pm$0.0 (14.3%) 0.22$\pm$0.0 (12.7%)
1 AR Equiv 1000 1.73$\pm$0.08 0.453$\pm$0.01 0.648$\pm$0.01 13.10$\pm$0.58 (1.0%) 0.20$\pm$0.0 (10.8%) 0.21$\pm$0.0 (10.6%)
1 AR Equiv 2500 1.41$\pm$0.01 0.513$\pm$0.01 0.633$\pm$0.03 11.76$\pm$0.91 (1.4%) 0.27$\pm$0.0 (28.8%) 0.27$\pm$0.0 (28.0%)
1 AR Equiv 5000 1.29$\pm$0.02 0.538$\pm$0.03 0.612$\pm$0.02 11.50$\pm$0.64 (1.9%) 0.28$\pm$0.0 (32.1%) 0.28$\pm$0.0 (26.2%)
1 AR, rand Equiv 5000 1.59$\pm$0.16 0.531$\pm$0.04 0.621$\pm$0.04 11.87$\pm$1.06 (1.9%) 0.26$\pm$0.0 (28.1%) 0.26$\pm$0.0 (24.1%)
1 AR Inv 5000 1.32$\pm$0.04 0.531$\pm$0.01 0.585$\pm$0.03 11.70$\pm$0.56 (1.3%) 0.26$\pm$0.0 (24.8%) 0.25$\pm$0.0 (20.1%)
1 NAR Inv 5000 1.54$\pm$0.04 0.571$\pm$0.00 0.430$\pm$0.02 14.26$\pm$0.51 (1.3%) 0.19$\pm$0.0 (15.9%) 0.18$\pm$0.0 (12.7%)
1 NAR Equiv 5000 1.46$\pm$0.06 0.584$\pm$0.00 0.473$\pm$0.02 13.04$\pm$0.88 (1.3%) 0.23$\pm$0.0 (24.0%) 0.22$\pm$0.0 (17.9%)
3 AR Equiv, DS 5000 1.23$\pm$0.05 0.539$\pm$0.01 0.620$\pm$0.01 11.47$\pm$1.05 (2.5%) 0.28$\pm$0.0 (31.4%) 0.28$\pm$0.0 (27.2%)
5 AR Equiv, DS 5000 1.25$\pm$0.01 0.539$\pm$0.02 0.596$\pm$0.03 11.90$\pm$1.00 (2.9%) 0.27$\pm$0.0 (31.6%) 0.26$\pm$0.0 (26.4%)
Groundtruth sequence prediction baseline: - 1.000$\pm$0.00 0.686$\pm$0.00 5.23$\pm$0.07 (27.9%) 0.56$\pm$0.0 (68.7%) 0.55$\pm$0.0 (68.7%)
Random sequence prediction baseline: - 0.251$\pm$0.00 0.012$\pm$0.00 24.40$\pm$0.34 (0.0%) 0.04$\pm$0.0 (0.0%) 0.02$\pm$0.0 (0.0%)
ViennaRNA 2D-only baseline: - 0.259$\pm$0.00 0.611$\pm$0.00 20.34$\pm$0.10 (0.0%) 0.07$\pm$0.0 (0.6%) 0.07$\pm$0.0 (1.1%)
Multi-state split 1 AR Equiv 5000 1.51$\pm$0.01 0.481$\pm$0.00 0.573$\pm$0.04 21.83$\pm$0.53 (0.0%) 0.12$\pm$0.0 (2.6%) 0.15$\pm$0.0 (5.5%)
3 AR Equiv, DS 500 1.87$\pm$0.04 0.444$\pm$0.01 0.587$\pm$0.02 22.09$\pm$0.13 (0.0%) 0.12$\pm$0.0 (2.3%) 0.14$\pm$0.0 (5.7%)
3 AR Equiv, DS 1000 1.76$\pm$0.04 0.455$\pm$0.03 0.504$\pm$0.04 22.92$\pm$1.43 (0.0%) 0.11$\pm$0.0 (2.3%) 0.14$\pm$0.0 (5.8%)
3 AR Equiv, DS 2500 1.54$\pm$0.07 0.500$\pm$0.01 0.543$\pm$0.01 22.00$\pm$0.26 (0.0%) 0.11$\pm$0.0 (2.9%) 0.14$\pm$0.0 (3.7%)
3 AR Equiv, DS 5000 1.44$\pm$0.04 0.531$\pm$0.00 0.573$\pm$0.03 22.19$\pm$0.28 (0.0%) 0.12$\pm$0.0 (4.2%) 0.15$\pm$0.0 (7.5%)
3 AR Equiv, DSS 5000 1.37$\pm$0.04 0.540$\pm$0.03 0.574$\pm$0.03 22.20$\pm$0.43 (0.0%) 0.12$\pm$0.0 (4.0%) 0.15$\pm$0.0 (7.5%)
5 AR Equiv, DS 5000 1.37$\pm$0.03 0.510$\pm$0.00 0.514$\pm$0.00 21.80$\pm$0.08 (0.0%) 0.12$\pm$0.0 (2.9%) 0.14$\pm$0.0 (6.2%)
1 NAR Equiv 5000 1.81$\pm$0.03 0.489$\pm$0.00 0.372$\pm$0.03 24.18$\pm$0.63 (0.0%) 0.09$\pm$0.0 (2.2%) 0.12$\pm$0.0 (4.7%)
3 NAR Equiv, DS 5000 1.65$\pm$0.13 0.506$\pm$0.01 0.346$\pm$0.02 24.06$\pm$0.43 (0.0%) 0.08$\pm$0.0 (2.0%) 0.11$\pm$0.0 (2.9%)
3 NAR Equiv, DSS 5000 1.60$\pm$0.10 0.520$\pm$0.02 0.352$\pm$0.03 24.18$\pm$0.55 (0.0%) 0.09$\pm$0.0 (2.2%) 0.12$\pm$0.0 (4.7%)
5 NAR Equiv, DS 5000 1.59$\pm$0.21 0.517$\pm$0.01 0.339$\pm$0.01 24.16$\pm$0.75 (0.0%) 0.08$\pm$0.0 (2.2%) 0.10$\pm$0.0 (4.5%)
Groundtruth sequence prediction baseline: - 1.000$\pm$0.00 0.525$\pm$0.00 17.52$\pm$0.32 (3.9%) 0.25$\pm$0.0 (24.2%) 0.29$\pm$0.0 (31.4%)
Random sequence prediction baseline: - 0.249$\pm$0.00 0.013$\pm$0.00 31.00$\pm$0.20 (0.0%) 0.03$\pm$0.0 (0.0%) 0.02$\pm$0.0 (0.0%)
ViennaRNA 2D-only baseline: - 0.258$\pm$0.00 0.470$\pm$0.00 29.10$\pm$0.00 (0.0%) 0.05$\pm$0.0 (0.0%) 0.05$\pm$0.0 (0.0%)
Split. Single- and multi-state splits are described in Section 3; the multi-state split is relatively harder than the single-state split based on overall reduced performance for all baselines and models. The multi-state split evaluates a particularly challenging o.o.d. scenario as the RNAs in the test set have significantly higher structural flexibility compared to those in the training set.
Max. #states We evaluate the impact of increasing the maximum number of states as input to gRNAde. Multi-state models improve native sequence recovery as well as structural self-consistency scores over an equivalent single state variant. Notably, on the more challenging multi-state split, the improvement in sequence recovery was observed to be as high as 5-6% for the best multi-state models. This trend holds even for the single-state benchmark where the multi-state model is being used with only one state as input. This suggests that seeing multiple states during training can be useful for teaching gRNAde about RNA conformational flexibility and improve performance even for single-state design tasks.
GNN and pooling architecture We ablated whether the internal representations of the GVP-GNN are rotation invariant or equivariant. Equivariant GNNs are theoretically more expressive (Joshi et al., 2023) and we find them more capable at fitting the training distribution (as shown by lower perplexity) which in turn results in improved metrics compared to invariant GNNs.
Model and decoder ‘AR’ implies autoregressive decoding (described in Section 2.2, uses 4 encoder and 4 decoder layers), while ‘NAR’ implies non-autoregressive, one-shot decoding using an MLP (uses 8 encoder layers). Across both evaluation splits, AR models show significantly higher self-consistency scores than NAR, even though NAR lead to higher sequence recovery for the single-state split. AR is more expressive and can condition predictions at each decoding step on past predictions, while one-shot NAR samples from independent probability distributions for each nucleotide. Thus, AR is a better inductive bias for predicting base pairing and base stacking interactions that are drivers of RNA structure (Vicens & Kieft, 2022). For instance, G-C and A-U pairs can often be swapped for one another, but non-autoregressive decoding does not capture such paired constraints.
Additionally, we also present results for the impact of training gRNAde with random decoding order. This can be practically very useful for partial or conditional design scenarios, and leads to a minor reduction in sequence recovery and 3D self-consistency (in line with what was observed for ProteinMPNN).
2D inverse folding baseline. We additionally report results for ViennaRNA’s 2D-only inverse folding method to further demonstrate the utility of 3D inverse folding. ViennaRNA has improved 2D self-consistency scores over gRNAde but fails to capture tertiary interactions in its designs, as evident by poor recovery and 3D self-consistency scores similar to the random baseline. We observed the same trend for other 2D-only inverse folding methods such as NuPack’s design tool. This result should not be surprising, as 2D tools are meant for design scenarios that only involve base pairing and do not take any 3D information into account.
Choice of structure predictors. As previously noted, self-consistency metrics are highly dependent on the performance of the structure prediction method used. We chose EternaFold as it is simple to use as well as validated for designed and synthetic RNAs, unlike most other 2D structure prediction tools. Replacing EternaFold with RNAFold lead to unchanged results and did not modify the relative rankings of the models:
  • AR, 1 state, Equiv. GNN, EternaFold scMCC: 0.612$\pm$0.02, RNAFold scMCC: 0.614$\pm$0.03.
  • NAR, 1 state, Equiv. GNN, EternaFold scMCC: 0.473$\pm$0.02, RNAFold scMCC: 0.477$\pm$0.04.
Lastly, we would like to note the challenge of evaluating multi-state design: Structural self-consistency metrics are not ideal for evaluating RNAs which do not have one fixed structure/undergo changes to their structure. It would be ideal (but extremely slow and expensive) to run MD simulations to validate multi-state design models.
Appendix C Additional Results
Table 2: Full results for Figure 4 comparing gRNAde to Rosetta, FARNA, ViennaRNA and RDesign for single-state design on 14 RNA structures of interest identified by Das et al. (2010). Rosetta and FARNA recovery values are taken from Das et al. (2010), Supplementary Table 2.
ViennaRNA FARNA RDesign Rosetta gRNAde (single-state)
PDB ID Description Recovery Recovery Recovery Recovery Recovery Perplexity 2D self-cons.
1CSL RRE high affinity site 0.25 0.20 0.4455 0.44 0.5719 1.2812 0.8644
1ET4 Vitamin B12 binding RNA aptamer 0.25 0.34 0.3929 0.44 0.6250 1.3457 -0.0135
1F27 Biotin-binding RNA pseudoknot 0.30 0.36 0.3013 0.37 0.3437 1.6203 0.4523
1L2X Viral RNA pseudoknot 0.24 0.45 0.3727 0.48 0.4721 1.3181 0.5692
1LNT RNA internal loop of SRP 0.33 0.27 0.5556 0.53 0.5843 1.4337 0.1379
1Q9A Sarcin/ricin domain from E.coli 23S rRNA 0.27 0.40 0.4417 0.41 0.5044 1.3411 0.0597
4FE5 Guanine riboswitch aptamer 0.29 0.28 0.4112 0.36 0.5300 1.3824 0.9116
1X9C All-RNA hairpin ribozyme 0.26 0.31 0.3967 0.50 0.5000 1.3905 0.6630
1XPE HIV-1 B RNA dimerization initiation site 0.27 0.24 0.3834 0.40 0.7037 1.2177 0.7768
2GCS Pre-cleavage state of glmS ribozyme 0.25 0.26 0.4518 0.44 0.5078 1.3053 0.4062
2GDI Thiamine pyrophosphate-specific riboswitch 0.25 0.38 0.3523 0.48 0.6500 1.2363 -0.0251
2OEU Junctionless hairpin ribozyme 0.23 0.30 0.5000 0.37 0.9519 1.0913 0.7768
2R8S Tetrahymena ribozyme P4-P6 domain 0.27 0.36 0.5641 0.53 0.5689 1.1881 0.7281
354D Loop E from E. coli 5S rRNA 0.28 0.35 0.4458 0.55 0.4410 1.4938 0.0430
Overall recovery: 0.27 0.32 0.4296 0.45 0.5682
Appendix D 3D Visualisation of gRNAde Designs
@@ Appendix C Additional Results
Figure 9: Cherry-picked designs for Guanine riboswitch aptamer (PDB: 4FE5, sequence: GGACAUAUAAUCGCGUGGAUAUGGCACGCAAGUUUCUACCGGGCACCGUAAAUGUCCGACUAUGUCC). We show the RhoFold-predicted 3D structure in colour overlaid on the groundtruth structure in grey. Designs recover the base pairing patterns and tertiary structure of the RNA, as measured by high self-consistency score. gRNAde’s perplexity is correlated well with 3D self-consistency scores and can be useful for ranking designs.
@@
GGGGCUCCGGCGACGCAGUCGA
AAGCCCAGCAGUACCAAGCCUC
AGGGGAAACUUUGAGGUGGCCU
AACAAAGGAUACGGUAAUAAGC
UGCGGGAAAAGGUUGUAAGCCG
GAGCGAAGACCUAAGGCACCGC
UUUUGGCGGUGCUAUGGUUGAA
GUUAA
@@
GGGGUACCGGCGACGCAGUCG
AAUGCCCUGUGGUACCAAGCC
CCGGGGGAAACUUCGGGGUGG
CCUUACCAAGGACACGGUAAU
AAGCCACGGGAAAUGGUUGUA
AGCCGGUCCGAAGCCCUAAGG
CCGCGCUUUGGGCGCGGCUAU
GGGUGAAGGCAA
@@
GAGGCCACGGCAACGCAGUC
UAACGCCCUGUGGUACCAAG
UCUUAGGAGAAAUUUUAAGA
UGGCCUAAUAAAGGAUAUGG
UAAUAAGCCACGGGAAAAGG
UUGUAAGACGUGACGAAGUC
CUAAGGCCACAGUUUUGCUG
UGGCUAUGGAUGGAGUACA
@@
Figure 10: Cherry-picked designs for Tetrahymena Ribozyme P4-P6 domain (PDB: 2R8S, sequence: GGAAUUGCGGGAAAGGGGUCAACAGCCGUUCAGUACCAAGUCUCAGGGGAAACUUUGAGAUGGC
CUUGCAAAGGGUAUGGUAAUAAGCUGACGGACAUGGUCCUAACACGCAGCCAAGUCCUAAGUCAACAGAUC
UUCUGUUGAUAUGGAUGCAGUUCA